Determine the benefit cost ratio for mango farm at rotation age of 7years given the following

1. At the rotation age, the expected yield is 221 cubic meter per hectare

2. Stumpage price is $100/m³

3. Total land area is 250 hectares

4. Plantation establishment cost which is assumed to be incurred at yield 2021 is $3,500 per hectare

5. Weeding cost at year 1 is $200 per hectare

6. Fertilizing cost incurred at year 2 is $200 per hectare per year

7. Interest rate is assumed to be 18%


1
Expert's answer
2021-05-04T07:38:25-0400

Determination o f Present Value of Cash Flows

1. Computation of Present Value of Cash Inflows (Benefit to be derived):

Mango Plantation has a rotational/cultivation age of 7 years. Therefore, the seeds sown now will reap benefits after 7 years. 

  • Per Hectare expected Yield: 221m3
  • Total Cultivation Area: 250 Hectares
  • Price to be derived: $100/m3

Total Cash Benefit to be derived after 7 Years:

"= Per" "Hectare" "Yield" "\times Total" "Cultivation" "Area \times" "Price" "to" "be" "derived"

"=221m^3\times 250" "hectares \times" $"100=" $"5,525,000"


PRESENT VALUE OF CASH INFLOW (BENEFIT):


"\frac{5,525,000}{(1+0.18)^7}=" $"1,734,436"


2. Computation of Present Value of Cash Outflow (Costs):



Computation of Benefit Cost Ratio


  • Present Value of Benefit: $5,525,000
  • Present Value of Cost: $1,107,949

Benefit Cost Ratio:

= Present Value of Cash Inflows (Benefit) / Present Value of Cash Outflows (Costs)

"=\frac{5,525,000}{1,107,949}=4.98"

BCR Ratio more than 1 shows that it is worth accepting the Project. i.e $1 of cost will yield a benefit of $4.98.